Suppose milk and honey are the only products produced in Ghana. Use the data belowfrom the
ii.real GDP, (3marks)
a.Computefor each year the
economy of Ghana to answer the questions that follow:
iii.the GDP deflator.(3marks)
YearPrice of MilkQuantity of MilkPrice of HoneyQuantity of Honey
i.nominal GDP, (3marks)
2012Ghc1200Ghc2100
2013Ghc2200Ghc4100
2011Ghc1100Ghc250
Base year: 2011
b.Compute for 2012 and 2013 from the respective preceding years the percentage change in
Required:
i.nominal GDP, (2marks)
ii.real GDP, (2marks)
iii.the GDP deflator (2marks)
1
Expert's answer
2021-08-02T11:05:42-0400

Solution:

a.). i.). Nominal GDP:

2011: (1 "\times"200) + (2 "\times"50) = 100 + 100 = 200

2012: (1 "\times"200) + (2 "\times"100) = 200 + 200 = 400

2013: (2 "\times"200) + (4 "\times"100) = 400 + 400 = 800

 

ii). Real GDP (Base year 2011):

2011: (1 "\times"200) + (2 "\times"50) = 100 + 100 = 200

2012: (1 "\times"200) + (2 "\times"100) = 200 + 200 = 400

2013: (1 "\times"200) + (2 "\times"100) = 200 + 200 = 400

 

iii.). GDP Deflator:

2011: "\frac{200}{200} \times100 = 100"


2012: "\frac{400}{400} \times100 = 100"


2013: "\frac{800}{400} \times100 = 200"


b.) i.). Percentage change in Nominal GDP:

2012: "(\frac{400 - 200}{200}) \times 100 = 100\%"


2013: "(\frac{800 - 400}{400}) \times 100 = 100\%"


ii.). Percentage change in Real GDP:

2012: "(\frac{400 - 200}{200}) \times 100 = 100\%"


2013: "(\frac{400 - 400}{400}) \times 100 = 0\%"

 

iii.). Percentage change in GDP Deflator:

2012: "(\frac{100 - 100}{100}) \times 100 = 0\%"


2013: "(\frac{200 - 100}{100}) \times 100 = 100\%"